### v2

# ICE-SEARCH: A Language Model-Driven Feature Selection Approach

Abstract
This study unveils the In-Context Evolutionary Search (ICE-SEARCH) method, the first work that melds language models (LMs) with evolutionary algorithms for feature selection (FS) tasks and demonstrates its effectiveness in Medical Predictive Analytics (MPA) applications. ICE-SEARCH harnesses the crossover and mutation capabilities inherent in LMs within an evolutionary framework, significantly improving FS through the model’s comprehensive world knowledge and its adaptability to a variety of roles. Our evaluation of this methodology spans three crucial MPA tasks: stroke, cardiovascular disease, and diabetes, where ICE-SEARCH outperforms traditional FS methods in pinpointing essential features for medical applications. ICE-SEARCH achieves State-of-the-Art (SOTA) performance in stroke prediction and diabetes prediction; the Decision-Randomized ICE-SEARCH ranks as SOTA in cardiovascular disease prediction. Our results not only demonstrate the efficacy of ICE-SEARCH in medical FS but also underscore the versatility, efficiency, and scalability of integrating LMs in FS tasks. The study emphasizes the critical role of incorporating domain-specific insights, illustrating ICE-SEARCH’s robustness, generalizability, and swift convergence. This opens avenues for further research into comprehensive and intricate FS landscapes, marking a significant stride in the application of artificial intelligence in medical predictive analytics.
1 Introduction
The proposed In-Context Evolutionary Search (ICE-SEARCH) method is the first work that melds anguage models (LMs) with evolutionary algorithms for feature selection (FS) tasks and demonstrates its effectiveness in Medical Predictive Analytics (MPA) applications.
Medical Predictive Analytics (MPA) represents a critical and specialized sector within health informatics dedicated to the sophisticated examination of medical data. MPA leverages advanced statistical algorithms and machine learning methodologies to forecast upcoming trends in healthcare outcomes, including patient admission rates, pharmaceutical advancements, and strategies for disease management (Van Calster et al. (2019)).
The core value of MPA resides in its ability to sift through and analyze the extensive datasets prevalent in the healthcare industry. This analytical power supports more informed clinical decision-making, enhances patient care quality, and facilitates more efficient allocation of healthcare resources. (Lamba et al. (2021)).
However, the sheer volume and complexity of this data pose significant challenges in converting it into actionable insights. In this context, the process of feature selection (FS) becomes critically important. By identifying the most pertinent subset of features within vast datasets, FS facilitates a more efficient analysis. Specifically, in medical scenarios like stroke prediction, where data is both extensive and varied, FS proves invaluable. It not only circumvents the "curse of dimensionality" but also aids in pinpointing disease causes. This, in turn, can unveil a new understanding of disease mechanisms, thereby contributing to advancements in medical research and patient care strategies (Remeseiro and Bolón-Canedo (2019)).
As highlighted in Section 2, FS confronts substantial hurdles, chiefly due to its broad search space that leads to an exponential increase in computational demands. Complications such as diminished generalizability, redundancy within intricate problem scenarios, challenges in handling high-dimensional data, an incomplete understanding of underlying data structures, and the obligation to presume data distributions that may not hold true, further exacerbate these challenges.
Contrasting this, the utility of Language Models (LMs) has significantly evolved, extending far beyond traditional text analysis. This evolution positions LMs as critical instruments in specialized fields. The adaptability and advanced capabilities of these models are demonstrated by their leading-edge performance in various sectors, including medical diagnostics Singhal et al. (2022), financial analysis Wu et al. (2023), data preprocessing Borisov et al. (2023), code generation Rozière et al. (2024) etc. This broad applicability underscores the potential of LMs to revolutionize not only how we approach complex datasets but also how we extract and interpret the most meaningful insights from them.
We posit that LMs can effectively contribute to FS in various applications. Furthermore, as suggested by Chen et al. (2023), semantic knowledge encoded in the LM allows for access to a considerable amount of cross-domain knowledge that would not be available in evolutionary algorithms, greater diversity of the outputs, and reduces the amount of manual design and human bias involved in the evolutionary algorithm, thereby powering evolutionary neuro-architecture search.
Hence, we introduce ICE-SEARCH, a FS methodology that integrates the in-context learning capabilities of language models Dong et al. (2023) with an evolutionary algorithmic framework. Our method involves engaging the LM in diverse roles, asking it to identify crucial features for a specific downstream task, and providing the LM with feedback on the performance of both effective and ineffective feature sets, measured in terms of training and validation accuracies of 10-fold cross-validation(CV). We also discover that during evolutionary in-context training, there is an enhancement in the model’s ability to uncover a set of high-quality feature combinations. At the time of submitting this manuscript, our research appears to be the first to explore the potential of LMs in this context.
To demonstrate the efficacy of our approach, we analyze the performance of ICE-SEARCH on three critical tasks–stroke prediction, cardiovascular disease prediction, and diabetes prediction–using three increasingly complex datasets. Each of these diseases is among the leading causes of morbidity and mortality worldwide, carrying substantial risks of severe health complications, long-term disability, and death.
We demonstrate that ICE-SEARCH consistently outperforms the existing FS methods in stroke prediction and diabetes prediction, and can be easily adapted to improve its performance on cardiovscular disease prediction whose dataset suffers from a discrepancy between training/valitaion and testing distribution. More interestingly, our empirical studies show that ICE-SEARCH is robust against poor population initialization and converges rapidly.
We summarize our contributions below:
  1. 1.
    ICE-SEARCH is the first work to apply an LM to dynamically search for valuable feature sets through role-playing and improve its in-context prompting examples without any fine-tuning.
  2. 2.
    ICE-SEARCH achieves SOTA, our experiments on stroke prediction and diabetes prediction demonstrate the effectiveness, robustness, stability, and scalability over existing approaches.
  3. 3.
    The value of ICE-SEARCH lies in its ability to empower end-users with cross-domain expertise encoded in a LM fine-tuned with domain-specific knowledge. ICE-SEARCH can easily be adapted to perform FS in areas like genetic analysis that commonly have thousands or even millions of features.
2 Related Work
Feature Selection: FS presents itself as a difficult task mainly due to two challenges: (i) its vast search space, where the total possible solutions are $2^{n}$ for a dataset with n features, leading to exponential growth in computing expenses, and (ii) the sub-optimal convergence caused by stagnations in local optima. Most of the techniques applied, such as heuristic search and random search, suffer from poor generalizability of evaluation criteria, redundant feature subsets when applied to complex problems Xue et al. (2016a), huge search space for high-dimensional data whose efficiency deteriorates quickly or computationally infeasible Miao and Niu (2016); Li et al. (2017), limited knowledge of the underlying data strucutre Li et al. (2017), or imposing case-specific assumption(s) about the probability of the distribution of the data Miao and Niu (2016). Despite that evolutionary computation(EC) methods have shown promises Xue et al. (2016a), flexible evaluation measures according to specific feature representations and search mechanisms that incorporate worldly knowledge across multiple disciplines are still in need for their optimal performance Li et al. (2017). Feature selection algorithms that are scalable, robust against dataset quality, and efficient are preferred Doquet and Sebag (2020); Li et al. (2017).
Evolutionary Algorithms for Feature Selection: Due to the vast search space, finding an optimal feature subset remains a challenging task. As discussed in the survey written by Chandrashekar and Sahin (2014), various ML and statistical FS algorithms have been proposed to tackle the task. In particular, as suggested by Xue et al. (2016b), EC-based approaches have recently been proven to be effective in many scenarios. Specifically, EC methods for feature selection have been applied in scenarios including genetic algorithm Holland (1992), particle swarm optimization Wang et al. (2018), genetic programming Muni et al. (2006), and correlation-based filter measures Hall (1999). Apart from the prospect of evolutionary algorithms for FS, it points out several issues and challenges that need to be addressed: scalability, computation cost, search mechanisms, stability, etc. Specifically, the FS algorithms need to: (1) have the capacity to process thousands or even millions of features, (2) have an efficient search strategy and a fast evaluation measure, (3) explore the whole search space and exploit the local regions if needed, and (4) be stable when using different starting points.
LLM for Feature Engineering: Feature engineering is closely related to feature selection, the former focuses on constructing new features from the original feature pool while the latter aims to identify and remove redundant and informative features. The article by Hollmann et al. (2023) introduces Context-Aware Automated Feature Engineering (CAAFE), a method leveraging LLMs for automated feature engineering in tabular datasets. CAAFE uses LLMs to iteratively generate both the Python codes for creating new semantically meaningful features and explanations for their utility. The method improved performance on 11 out of 14 datasets, demonstrating its effectiveness over traditional methods. Similarly, CAAFE also combines the domain knowledge encapsulated in LLMs with classical machine learning algorithms, offering a novel approach to integrating domain knowledge into automated machine learning processes.
Evolutionary Prompting: Prompting is a technique employed to give LLMs leverage to address specialized tasks more efficiently. However, the performance of LLMs varies contingent upon the selection of the prompt, and that leads to the topic of automatic prompt optimizations, such as continuous prompt tuning and discrete prompts Liu et al. (2023); Li and Liang (2021); Zhang et al. (2021). In contrast, the high variance in prompt effectiveness for downstream tasks has led to the development of prompt generation, prompt revision, and prompt editing methods that focus on the exploration or exploitation of candidate prompts to select the most optimal one(s) Zhou et al. (2022); Guo et al. (2023a); Zhang et al. (2022). To address the constraints of prompts designed for predefined answers, which may encounter challenges when applied to generation tasks, and to mitigate the risk of encountering local optima, recently introduced evolutionary prompting techniques offer the advantages of not relying on predefined parameters or gradients while simultaneously maintaining a balance between exploration and exploitation Guo et al. (2023b); Chen et al. (2023). Such paradigms of evolutionary prompting inspired ICE-Search, which adapts and evolves in the search for optimal prompts to identify critical features for downstream tasks. Similar to evolutionary prompting, ICE-Search showcases novel ways of leveraging language models while offering flexibility and adaptability to diverse metrics and tasks.
3 ICE-SEARCH: Feature Selection with Language Model
Via zero-shot prompting and in-context learning, ICE-SEARCH aims to streamline the evolutionary training approach for FS in MPA. ICE-SEARCH utilizes an LM as the crossover and mutation operator and directs it to dynamically identify important feature collections through simulating medical practitioners’ roles, thus enriching its in-context prompting examples without needing fine-tuning.
As evidenced by section 4.5, the identified features are indicative of the risk factors of diseases and can improve the accuracy of MPA. ICE-Search’s ability to extract representative features from medical datasets emphasizes its crucial role in MPA. This capability allows healthcare professionals to develop precise and efficient predictive models, enabling the optimization of healthcare resources for early disease detection. In this section we explain our algorithm in detail.
Notations
  • $\mathcal{X}$: A mapping that uses a set of X number of classical FS methods used to initialize the feature set pool. Output X feature sets.
  • Y: Number of times for the LM to select feature sets through zero-shot prompting for feature set pool initializaion.
  • $\mathcal{Z}$: Set of role-plays for the LM intended for evolution.
  • U: Number of top feature sets to keep.
  • V: Number of bottom feature sets to keep.
  • $N$: Number of groups that the given dataset is split into for CV.
  • $E$: Number of Epochs.
  • $\mathcal{F}$: The given feature set
  • $\mathcal{S}$: Feature set pool (i.e. population)
  • $\mathcal{D}$: Dataset.
  • $\mathcal{T}$: Task description.
  • $\Phi(\mathcal{D},s,N)$: Evaluation of the input feature set $f$ usig a pre-determined downstream model on dataset $\mathcal{D}$, output average training and testing accuracies across N-fold CV.
[…]
Figure 1: An illustration of ICE-SEARCH Architecture
Algorithm 1 In-Context Evolutionary Search (ICE-SEARCH)
1:hyperparameters: $\mathit{Y}$: int, $\mathit{U}$: int, $\mathit{V}$: int, N: int, E: int
2:input: $\mathcal{D}$, $\mathcal{F}$, $\mathcal{T}$, $\mathcal{Z}$, $\mathcal{X}(\cdot)$, $\mathcal{P}(\cdot)$, $\Psi(\cdot)$, $\Phi(\cdot)$,
3:procedure
4:     $\mathcal{S}\leftarrow\mathcal{X}(\mathcal{F})$
5:     $\mathcal{S}[s]\leftarrow\Phi(\mathcal{D},s,N)$ for $s$ in $\mathcal{S}$
6:     $P\leftarrow\mathcal{P}(\mathcal{T},\mathcal{F})$
7:     TEMPSET = {}
8:     for $i\leftarrow 1$ to $\mathit{Y}$ do
9:         $s\leftarrow\Psi(P,\text{ROLE})$
10:         train acc, val acc = $\Phi(\mathcal{D},s,N)$
11:         $\texttt{TEMPSET}[s]\leftarrow$ train acc, val acc
12:     end for
13:     $S\text{.append}({\mathrm{argmax}_{s}}\ \texttt{TEMPSET}[s][1])$
14:     for $e\leftarrow 1$ to E do
15:         P = $\mathcal{P}(\mathcal{T},\mathcal{F},S)$
16:         for $z$ in $\mathcal{Z}$ do
17:
$f\leftarrow\Psi(P,z)$
18:              train acc, val acc = $\Phi(\mathcal{D},s,N)$
19:              $\mathcal{S}[s]\leftarrow$ train acc, val acc
20:         end for
21:         Sort $\mathcal{S}$ based on validation accuracy
22:         $\mathcal{S}\leftarrow$ $\mathcal{S}$[:U] $\bigcup$ $\mathcal{S}$[-V:]
23:     end for
24:     $s\leftarrow$ feature set in $\mathcal{S}$ that achieves the highest validation accuracy evaluated by $\Phi$
25:     if ties exist then
26:
choose among the tied feature sets that have the lowest training accuracy
27:     end if
28:     return $s$
29:end procedure
3.1 Algorithm Archetecture
ICE-SEARCH architecture can be broken down into four components: Initialization, Crossover and Mutation, Filtration, Final Selection. Figure 1 provides a high-level overview of our algorithm and Algorithm 1 illustrates its low-level implementation.
Initialization (line 4-12): The first stage of ICE-SEARCH is the initialization of a population $\mathcal{S}$, which is a pool of feature sets. This initial population is derived using $\mathit{X}$ number of classical FS methods, with each method’s performance assessed based on average training and validation accuracies of the downstream model $\Phi$ over N-fold CV. To further refine the initial population, we prompt a LM through $\Phi(\cdot)$ $Y$ times in a zero-shot manner, and the output feature combination that yields the highest average validation accuracy over N-fold CV is chosen and put in $\mathcal{S}$ for further processing.
Crossover and Mutation (line 13-20): In view of the achievements attributed to EVO Prompting by Chen et al. (2023) and EVO Prompt by Guo et al. (2023b), our evolutionary strategy leverages a LM as the mutation and crossover operator throughout each epoch. Specifically, the LM assumes $\mathcal{Z}$ distinct roles and is informed about the average training and validation accuracies associated with each feature set in $\mathcal{S}$. During each of these role-plays $\Psi$, the LM selects a set of features, and its average training and validation accuracies over N-fold CV is computed via $\Phi(\cdot)$ and recorded.
Filtration (line 21-22): After each epoch, we apply a filtering process. This entails retaining only the feature combinations that fall into two categories: those with the top $\mathit{U}$ highest validation accuracies and those with the bottom $\mathit{V}$ validation accuracies assessed via $\Phi(\cdot)$. This selective mechanism ensures that our algorithm exploits the most promising feature combinations’ potential while also maintaining a degree of exploration in $\mathcal{P}$.
Final Selection (line 24-28): The iterative evolutionary process is repeated for E epochs. Upon completion of the final epoch, we select the feature set that demonstrates the highest validation accuracy. In cases where multiple feature combinations exhibit the same highest validation accuracy, the tie-breaker is the combination with the lowest training accuracy. This approach ensures that our model not only performs well on unseen data but also avoids overfitting to the training dataset. On the other hand, as shown in table 5 and 8, we have found empirically that all $\mathit{U}$ feature sets are valuable as many of them yield high test accuracy. Therefore, it is not mandatory to follow strictly the strategy we just described. One can customize this component by adjusting step 24 to 27 in algorithm 1 according to the underlying task complexity or dataset quality.
3.2 Evolution-driven Prompting
ICE-SEARCH implements a unique and practical evolution-driven prompting mechanism depicted in Figure 2. This approach allows the LM to iteratively generate and evaluate new feature sets, thereby augmenting its own prompting capabilities. This multi-objective functionality offers several key benefits:
Figure 2: An illustration of ICE-SEARCH Promptings
Dynamic Learning and Adaptation: By continually generating and assessing new feature sets while eliminating old feature sets, the backbone LM evolves its understanding of what constitutes an effective feature combination for a given task via in-context learning. This evolutionary learning mechanism ensures that the model remains relevant and creates a continuous improvement loop that augments its prompt over time. Moreover,
Role-played Crossover and Mutation with LMs: A suitably prompted and sampled LM can provide rigorous insights to users’ questions, especially through role-playing the characters of facts-based AI assistants Shanahan et al. (2023). This capability stems from models’ exposure to factual examples and scientifically precise guidances within its training processes Shanahan et al. (2023). Consequently, the agent excels in role-playing as a reliable source of information, drawing on the factual correctness of statistics-driven and fact-based examples in contexts. Hence, ICE Search leverages an LM as the crossover and mutation operator by engaging it with various roles. By integrating LMs directly into the evolutionary search process, we effectively enhance the metrics of interest. This approach allows us to exploit the rich semantic information encoded in LMs, leading to more informed feature selection decisions.
Customization and Flexibility: The adaptability of the algorithm allows for a high degree of customization and flexibility. For example:
  • Custom Feature Sets for Different Domains: In medical datasets, the algorithm can be asked to prioritize features like patient history and laboratory results, whereas in financial datasets, it might focus on market trends and economic indicators.
  • Adaptabililty for Diverse Applications: The downstream algorithm, population initialization methods, number of epochs, feature pool size, and the roles assumed by the language model (LM) can all be tailored based on the task’s complexity and specific requirements. Additional parameters, such as early stopping criteria and maximum/minimum number of features, can also be integrated.
  • Flexible Evolution Strategies: Depending on the goal, the algorithm can be configured to perform exploitation and exploration based on different policies and metrics. For instance, one might consider including randomly chosen feature sets in the first few epochs.
  • Scalability to Different Dataset Sizes: Depending on the context length of the backbone LM, our algorithm can be adjusted to handle large-scale datasets that feature thousands or millions of variables, or it can be fine-tuned for smaller datasets to focus on detailed feature analysis.
3.3 Feature Selection and World Knowledge:
ICE-SEARCH leverages the LM’s extensive world knowledge and cross-domain expertise to facilitate the transfer of insights across related fields. For example, in the context of cardiovascular disease prediction, the algorithm can draw valuable insights from stroke prediction due to its clinical and biological characteristics. Leveraging knowledge from these related fields enhances the feature selection process, potentially revealing common or complementary predictors that hold significance across various health conditions. Using the vast information stored in an LM, ICE-SEARCH looks beyond a dataset and helps end-users identify significant features through analysis of interconnected domains.
4 Experiments and Evaluation
In this section, we demonstrate ICE-SEARCH’s effectiveness for FS in three tasks: stroke prediction, cardiovascular disease prediction, and diabetes prediction.
We begin by selecting four existing FS methods for initializing the population. The outputs of the four methods then interact with a language model evolutionarily. Note that no finetuning is required in our algorithm, we also examine the algorithm’s generalizability and robustness from experimental results.
4.1 Public Health Impact and Significance
The National Institute of Diabetes and Digestive and Kidney Diseases actively conducts and supports clinical trials targeting diseases with significant public health impact, including hypertensive heart disease (HHD). The global prevalence of HHD and absolute rates of adverse outcomes are expected to continue to rise due to population growth, aging, and continuing increases in obesity globally (Roth et al. (2020).
Fig 3 demonstrates the heightened global and regional impact of HHD, quantified by the Disability-Adjusted Life Year (DALY) per 100,000 persons — one DALY represents the loss of the equivalent of one year of full health (WHO ). The global DALYs attributable to HHD is 2150.8 with a mean of 171.93 and a standard deviation of 267.64 in 2019. They are notably elevated in Asia where DALYs range between 975.46 and 1417.03 (Yang et al. (2023).
The observed geographic disparity in DALYs due to HHD highlights the substantial global health burden it represents, particularly in the hardest-hit regions. This disparity underscores the critical need for effective strategies to identify the causes of HHD for both diagnosis and treatment. Current research is exploring a variety of areas, including risk factors specific to certain demographics, advanced diagnostic techniques, and the genetic basis of these conditions. (NIH ). Hence, any improvement on HHD MPA tasks would save a lot of lives.
In this study, we select three open-source datasets on stroke, cardiovascular disease, and diabetes, all of which are either direct complications or contributing factors to HHD, due to the systemic effects of prolonged hypertension on cardiovascular health (Petrie et al. (2018). These case studies are intended to demonstrate the efficacy of ICE-SEARCH in improving the MPA of diseases with high risk.
Figure 3: Geometric Patterns of DALYs Due to HHD in 2019. (Source of Data: Yang et al. (2023))
4.2 Datasets
Stroke: The stroke prediction dataset from Kaggle ( Fedesoriano (2021)) comprises 10 parameters including patient age, gender, medical history (specifically hypertension and heart disease), marital status, occupation type, residential environment, average glucose level, body mass index, and smoking status. Initially, an imbalanced distribution was observed within the dataset, characterized by a preponderance of negative cases (4861 instances) juxtaposed with positive stroke instances (249 instances). To correct this distributional disparity, we implemented the Synthetic Minority Over-sampling Technique (SMOTE) ( Chawla et al. (2002)), which equalizes the representation by creating 4861 cases for stroke and non-stroke instances respectively. Furthermore, to maintain the dataset’s integrity while preserving its inherent distribution, we employed a method of imputation utilizing the median values of the corresponding columns.
Cardiovascular Disease: The cardiovascular health dataset ( Ulianova (2019)) has a near-equal distribution of both outcomes. It contains 31,783 negative samples and 30,962 positive samples, and includes 11 features including age (in days), height (in centimetres), weight (in kilograms), gender (categorical code), systolic blood pressure, diastolic blood pressure (both in mmHg), cholesterol levels (categorized as normal, above normal, well above normal), glucose levels (similarly categorized), along with binary indicators for smoking, alcohol intake, and physical activity, with no missing values. The primary variable of interest is the presence or absence of cardiovascular disease, represented as a binary target.
Diabetes: The diabetes dataset ( Teboul (2021)) features 218K negative samples, 35K positive samples, and no missing entries. Similarly, to address the class imbalance issue, we applied the SMOTE which results in a balanced dataset with 218,334 negative and 218,334 positive samples. The dataset features 21 indicative variables: high blood pressure, high cholesterol, undergone cholesterol checks, body mass index, smoking status, history of stroke, history of heart disease or attack, engagement in physical activity, regular consumption of fruits, regular consumption of vegetables, heavy alcohol consumption, access to healthcare services, avoiding doctor visits due to cost, general health condition, mental health condition, physical health condition, difficulty in walking, gender, age, education level, income. The target variable has 2 classes, no diabetes and prediabetes/diabetes, represented by 0 and 1 respectively.
4.3 Experiment Design
In the experiment, we use Phi2 ( Javaheripi and Bubeck (2023)) and 4-bit quantized version of LLaMA2 7B( Touvron et al. (2023)) models respectively as the backbone LM, each containing fewer than 8 billion parameters. The rationale for selecting these LMs lies in their constrained parameter count which allows us to explore the lower boundaries of the performance of ICE-SEARCH.
Initialization: The initialization stage begins by incorporating the backbone LM and four FS algorithms: Decision Tree, Random Forest, Logistic Regression, and Fisher Score Grabczewski and Jankowski (2005); Jaiswal and Samikannu (2017); Cheng et al. (2006); Gu et al. (2012). The LM is subsequently asked to assume the role of a medical doctor to recommend a list of important features based on the given task and dataset features in a zero-shot manner. We prompted the LM five times and the best feature set selected is combined with those selected by the other four FS methods. This combination forms the initial population for our evolution algorithm.
Evolution: The evolution stage comprises 17 distinct roleplays chosen according to the downstream task: Neurologist, Cardiologist, Radiologist, Epidemiologist, Public Health Professional, Pharmacist, Genetic Counselor, Health Informatics Specialist, Data Scientist, Data Analyst, Machine Learning Engineer, Biostatistician, AI/ML Researcher, Data Engineer, Ethical AI Advocate, Nurse, Emergency Medicine Physician. The goal of these roleplays is to facilitate crossover and mutation operations within the language models. Note that we did not conduct any prompt optimization or search for the "ideal" roleplays.
Filtration: During filtration, we use a predetermined downstream prediction model to evaluate average test and validation accuracies over N-fold CV of each of the 17 outputs and concatenate them with the feature sets pool. The feature sets pool undergoes an ordering process, whereby only the first five and last three feature sets are retained. At the end of each epoch, the pool contains five good and three poor feature sets.
Downstream Model: we employ XGBoost ( Chen and Guestrin (2016)) and Support Vector Machine (SVM) as the downstream models and validated the accuracy of the feature sets through a 10-fold CV to mitigate risks of overfitting. We choose not to conduct an extensive hyperparameter search to emphasize the simplicity and practicality of ICE-SEARCH. The experiments are intended to demonstrate its effectiveness, robustness, generalizability, and potential, rather than claiming it as the state-of-the-art methodology for feature selection.
Experiment Control: The experiments are conducted over eight epochs and five seeds(42-46). Key adjustments in the language model parameters include setting the temperature to 1 and the top p-value to 0.9 to introduce controlled randomness into ICE-SEARCH.
4.4 Prompt Design
The algorithm adopts a two-step approach, encompassing zero-shot prompting and few-shot prompting. Note that no optimization is conducted for the prompt design.
Zero-Shot Prompting: During initialization, we present scenarios to the model without any prior training on similar tasks via Zero-Shot prompting. For instance, we use prompts like: “Imagine you are a medical doctor. I need you to recommend important features for accurately [TASK DESCRIPTION]. Consider the following features: [FEATURES]. Think step by step. Selected features: …”. The model generate responses by relying solely on its pre-existing knowledge.
Few-Shot Prompting: During the crossover and mutation phases of the algorithm, we adopt a more targeted approach via few-shot prompting. We provide samples of feature sets that had achieved the top five and bottom three validation accuracies in previous assessments. An example prompt for this phase is: “As a Neurologist, recommend important features for [TASK DESCRIPTION]. Consider the following features: [FEATURES]. Here are the selected features and their corresponding classification accuracy results: [DETAILED FEATURE AND ACCURACY LIST]. Be innovative and think step by step. Features selected: …”
Both zero-shot and few-shot prompts, as demonstrated in Fig 2, are crafted to not only direct the model’s attention to specific tasks but also to stimulate explorative and rational thinking.
4.5 Evaluation Results
To evaluate the effectiveness of ICE-SEARCH, we rank the testing and average validation accuracies for each possible feature combination in predicting stroke and cardiovascular diseases. Specifically, the stroke dataset includes 1,023 feature combinations, while the cardiovascular disease dataset includes 2,047 combinations. We calculate the validation accuracies using a 10-fold CV approach and record the average values. However, we do not perform ranking analyses for feature sets in the diabetes dataset. The reason is that assessing test and validation accuracies for each of the $2^{21}$ feature combinations becomes impractical for our study.
4.5.1 Stroke Prediction:
Population Initialization Metric Decision Tree Random Forest Logistic Fisher Score Ensemble of FS ICE-SEARCH
FS methods + Phi2 Test Rank 149.2 48.8 77.4 93.0 51.4 36.4
Val Rank 96.8 8.0 61.6 82.4 2.8 1.0
Worst Test Rank 963.8
Worst Val Rank 977.0
Poor Feature Sets + Phi2 Test Rank 38.2
Val Rank 1.6
Table 1: Average Ranks (out of 1023) of FS Algorithms on Stroke dataset, using XGBoost as Downstream Classifier and Phi2 as the Backbone LM of ICE-SEARCH.
Seed Decision Tree Random Forest Logistic Fisher Score ICES $1^{st}$ ICES $2^{nd}$ ICES $3^{rd}$ ICES $4^{th}$ ICES $5^{th}$
42 87.969, 23 88.072, 21 87.506, 65 86.221, 333 88.483, 6 88.38, 11 87.969, 23 87.506, 65 87.969, 23
43 85.758, 418 87.198, 69 86.889, 128 86.941, 112 86.889, 128 86.941, 112 86.581, 204 87.558, 27 87.095, 84
44 88.175, 37 88.38, 21 88.843, 4 88.997, 1 88.997, 1 88.689, 8 88.38, 21 87.918, 68 88.123, 43
45 87.301, 20 87.301, 20 86.838, 43 87.661, 8 87.661, 8 88.278, 5 88.535, 2 87.301, 20 87.712, 6
46 86.838, 248 87.455, 113 87.301, 147 88.329, 11 87.969,39 87.712, 73 87.455, 113 87.969, 39 87.969, 39
Table 2: Test Accuracies(%) of FS Algorithms and their ranks (out of 1023) on Stroke dataset, using XGBoost as Downstream classifier and Phi2 as the Backbone LM of ICE-SEARCH. ICES $n^{th}$ stands for the feature set with the $n^{th}$ highest validation accuracy chosen by ICE-SEARCH.
Seed Train Accuracy Validation Accuracy Test Accuracy
42 71.174 69.436 68.792
68.314 65.475 65.810
81.358 79.349 81.028
64.758 63.893 63.753
43 70.966 68.908 69.717
68.353 65.976 66.324
81.910 80.017 79.640
64.579 63.945 64.576
44 71.271 69.422 67.969
68.243 65.334 66.787
81.658 79.786 80.154
64.675 64.241 64.216
45 71.053 69.436 69.409
68.363 65.848 65.964
81.498 79.606 80.668
64.644 63.958 64.165
46 70.871 68.638 70.026
68.141 65.436 66.530
81.525 79.530 80.771
64.279 64.279 64.833
Table 3: The accuracy of the four suboptimal feature sets, selected to evaluate the robustness of ICE-SEARCH against poor initial feature set selection methods.
Metric Decision Tree Random Forest Logistic Fisher Score ICE-SEARCH
Test Rank 58.0 7.8 171.2 63.4 13.8
Val Rank 38.8 7.0 201.2 52.0 1.4
Worst Test Rank 781.2
Worst Val Rank 835.8
Table 4: Average Ranks (out of 1023) of FS Algorithms on Stroke dataset, using Support Vector Machine(SVM) as Downstream Classifier and Phi2 as the Backbone LM of ICE-SEARCH.
Seed Decision Tree Random Forest Logistic Fisher Score ICES $1^{st}$ ICES $2^{nd}$ ICES $3^{rd}$ ICES $4^{th}$ ICES $5^{th}$
42 80.566, 2 80.566, 2 79.743, 218 80.0, 138 80.617, 1 80.566, 2 80.566, 2 80.566, 2 80.411, 13
43 78.766, 246 79.229, 27 78.766, 246 79.126, 63 79.126, 63 79.126, 63 79.229, 27 79.075, 97 79.075, 97
44 80.36, 33 80.566, 1 80.463, 13 80.566, 1 80.566, 1 80.566, 1 80.566, 1 80.566, 1 80.514, 7
45 79.383, 7 79.383, 7 79.126, 128 79.383, 7 79.434, 2 79.383, 7 79.383, 7 79.383, 7 79.383, 7
46 79.846, 2 79.846, 2 79.229, 251 79.537, 108 79.846, 2 79.794, 12 79.846, 2 79.846, 2 79.846, 2
Table 5: Test Accuracies(%) of FS Algorithms and their ranks (out of 1023) on Stroke dataset, using SVM as Downstream classifier and Phi2 as the Backbone LM of ICE-SEARCH. ICES $n^{th}$ stands for the feature set with the $n^{th}$ highest validation accuracy chosen by ICE-SEARCH.
Metric Phi2 4bit Quantized LLaMA2 7B
Test Rank 36.4 36.4
Val Rank 1.0 1.0
Worst Test Rank 963.88 1022.6
Worst Val Rank 977.0 1023.0
Table 6: Average Ranks (out of 1023) of ICE-SEARCH with Different Backbone LMs on Stroke dataset, using XGBoost as Downstream Classifier.
Table 1 showcases the average rankings for both test and validation accuracies in the context of stroke prediction. These rankings are based on five trials for feature sets chosen by several methods: Decision Tree, Random Forest, Logistic Regression, Fisher Score, and ICE-SEARCH. ICE-SEARCH outperforms the other four feature selection methods with an average test ranking of 36.4 and a validation ranking of 1. This performance not only exceeds that of the individual methods but also surpasses the approach of employing the FS method among the four that produces the feature set with the highest validation accuracy, referred to as the Ensemble of FS strategy.
Interstingly, ICE-SEARCH also showcases its efficacy in identifying less predictive feature sets, with its test and validation accuracies averaging rankings of 963.8 and 977.0, respectively. This highlights ICE-SEARCH’s dual capability in distinguishing between effective and ineffective feature sets. This demonstrates its capacity to identify poor feature sets from a baseline of effective ones, thereby implying its effectiveness in uncovering high-quality feature sets even when starting with a low-quality initial population.
To test this hypothesis, we replace the feature sets chosen by the four methods in the initial pool with four sets of lower accuracy. The training, validation, and testing accuracies of the four methods are provided in Table 3. As shown in Table 1, these suboptimal feature sets achieve average validation and test accuracy rankings of 38.2 and 1.6 respectively. The results indicate ICE-SEARCH’s ability to target effect feature sets from a baseline of ineffective ones, thus underpinning ICE-SEARCH’s robustness against poor initial feature set selection methods.
Table 2 presents the test accuracies and their corresponding rankings for four FS methods and ICE-SEARCH. We note that ICE-SEARCH ranks first in three out of five seeds, outperforming/matching the other four FS methods. Remarkably, within the results, ICE-SEARCH’s $4^{th}$ and $3^{rd}$ iterations are highlighted as the superior feature sets in seed 43 and 45, respectively. This suggests that not only is ICE-SEARCH $1^{st}$ highly effective, but all top five feature sets identified through the ICE-SEARCH algorithm hold significant value.
Tables 45, and 6 detail the performance of ICE-SEARCH in stroke prediction, utilizing the Support Vector Machine (SVM) as the downstream classifier and the 4-bit quantized LLaMA 2 7B as the underlying LM. According to Table 4, although the average test ranking of ICE-SEARCH is 13.8, which is six places lower than the highest average test rank of 7.8, ICE-SEARCH consistently excels in identifying feature sets with the best validation accuracies, achieving an average ranking of 1.4. This performance illustrates ICE-SEARCH’s effectiveness not only in selecting effective feature sets but also in recognizing inadequate ones, as evidenced by the significant deterioration in the worst test and validation ranks from 171.2 and 201.2 to 781.2 and 835.8, respectively—an improvement in identifying less effective feature sets by about 110
Table 5 further demonstrates ICE-SEARCH’s ability to identify the optimal feature sets in four out of five trials. Impressively, it also efficiently discovers additional feature sets that exhibit top test accuracies. Specifically, in seed 42, ICE-SEARCH uncovers one more feature set with the second-highest test accuracy; in seed 46, it finds two additional feature sets also ranking second in test accuracy. This efficiency of ICE-SEARCH in uncovering multiple high-performing feature sets significantly aids researchers in the field of feature selection, offering a broader range of options for achieving optimal test accuracies.
To evaluate the ICE-SEARCH algorithm’s adaptability with different underlying Language Models (LMs), its performance was analyzed using a 4-bit quantized LLaMA2 7B. Table 6 offers a comparative analysis of the ICE-SEARCH feature selection (FS) mechanism when applied to Phi2 and LLaMA 2 7B models. Notably, both models achieved identical rankings for test and validation accuracies. However, the 4-bit quantized LLaMA2 7B demonstrated superior capability in identifying less effective feature sets. Specifically, it accurately identified the feature set with the lowest test and validation accuracies, which were ranked at 1022.6 and 1023, respectively.
4.5.2 Cardiovascular Disease:
Metric Decision Tree Random Forest Logistic Fisher Score Ensemble of Staitstics ICE-SEARCH
Test Rank 34.4 189.8 293.2 53.0 40.0 44.6
Val Rank 22.6 202.0 260.4 11.4 4.2 2.8
Worst Test Rank 1750.2
Worst Val Rank 1761.2
Table 7: Average Ranks (out of 2047) of FS Algorithms on Cardiovascular Disease dataset, using XGBoost as Downstream Classifier and Phi2 as the Backbone LM of ICE-SEARCH.
Seed Decision Tree Random Forest Logistic Fisher Score ICES $1^{st}$ ICES $2^{nd}$ ICES $3^{rd}$ ICES $4^{th}$ ICES $5^{th}$
42 73.907, 44 73.771, 86 73.257, 265 73.743, 98 73.907, 44 73.743, 98 73.943, 24 73.971, 19 73.729, 107
43 73.007, 92 73.05, 76 72.586, 316 72.95, 120 73.007, 92 72.95, 120 73.121, 40 73.079, 64 72.921, 137
44 73.279, 1 73.25, 5 72.471, 283 73.157, 29 73.086, 64 73.157, 29 73.107, 51 73.279, 1 73.179, 20
45 73.521, 13 72.0, 760 72.686, 270 73.507, 17 73.636, 1 73.521, 13 73.329, 93 73.5, 19 73.479, 24
46 74.193, 22 74.193, 22 73.207, 332 74.286, 1 74.193, 22 74.057, 85 74.2, 18 74.207, 13 74.286, 1
Table 8: Test Accuracies(%) of FS Algorithms and their ranks on Cardiovasclar Disease dataset, using XGBoost as Downstream classifier and Phi2 as the Backbone LM of ICE-SEARCH. ICES $n^{th}$ stands for the feature set with the $n^{th}$ highest validation accuracy chosen by ICE-SEARCH.
Table 7 and 8 illustrate the performance of ICE-SEARCH in terms of accuracy and their respective rankings in the cardiovascular disease dataset. In predicting cardiovascular disease, ICE-SEARCH showcases a strong performance with average rankings of 44.6 for test accuracies and 2.8 for validation accuracies. Despite its average testing accuracy ranking second to the four FS methods, ICE-SEARCH ranks first in validation accuracy. This highlights a mismatch between the distributions of the training and testing datasets, suggesting that selecting feature sets based solely on the highest validation accuracy may not guarantee high test accuracy for this dataset. For example, in Table 8, with the seed set at 42, ICE-SEARCH $3^{rd}$ and $4^{th}$ rank at 24th and 19th, outperforming selections from the Decision Tree and ICE-SEARCH $1^{st}$. Similarly, with seed 43, ICE-SEARCH $3^{rd}$ and $4^{th}$ rank at 40th and 64th respectively in terms of test accuracy, exceeding the performance of selections by Random Forest and ICE-SEARCH $1^{st}$. Furthermore, the lowest test and validation accuracies are ranked at 1750.2 and 1761.2 respectively, emphasizing ICE-SEARCH’s consistent ability to identify less significant features.
[…]
96 20.0 56.2
92 40.2 48.4
88 39.4 65.2
84 29.4 31.4
80 29.0 41.6
76 24.0 56.2
Table 9: Average Test Ranks (out of 2047) of Decision-Randomized ICE-SEARCH on Cardiovascular Disease dataset, using XGBoost as Downstream Classifier and Phi2 as the Backbone LM of ICE-SEARCH. Seven seeds were chosen as a means to maintain the reproducibility of the randomized decisions.
Given the discrepancy between training, validation and testing distributions, it may be beneficial to adjust steps  24-28 in Algorithm 1 to increase its variance. As a solution, we introduce an adapted version called Decision-Randomized ICE-SEARCH, which involves randomly selecting one feature set from the final pool rather than choosing the one with the highest validation accuracy.
Table 9 shows the performance of Decision-Randomized ICE-SEARCH. For enhanced reproducibility and reliability, we utilize seven seeds for the random selection process. The findings reveal that in six out of seven cases, Decision-Randomized ICE-SEARCH recommends feature sets with an average test accuracy rank lower than those selected by the standard ICE-SEARCH. Notably, in four out of six cases, it identifies feature sets whose performance surpasses those chosen by the four FS methods. This indicates ICE-SEARCH’s flexibility to adapt to mismatches between training/validation and testing dataset distributions.
Decision-Randomized ICES, an approach which excludes the first choice of ICE-SEARCH (ICE-SEARCH $1^{st}$) and randomly selects from the subsequent options, demonstrates that in five out of seven instances, it results in less than ideal outcomes. This finding implies that the random selection process is more crucial and that additional layers of selection or filtering may be unnecessary. The experiments underscore the significance of each of the five surviving feature sets, marking the final pool itself as a valuable resource.
4.5.3 Diabetes Prediction:
Seed Decision Tree Random Forest Logistic Fisher Score ICES $1^{st}$ ICES $2^{nd}$ ICES $3^{rd}$ ICES $4^{th}$ ICES $5^{th}$
42 88.552 90.931 90.613 89.713 90.959 91.002 90.927 90.931 90.897
43 91.017 89.627 90.694 90.752 90.987 91.034 90.992 91.017 90.953
44 90.924 90.87 89.721 90.566 90.93 90.906 90.934 90.87 90.924
45 88.244 90.77 90.715 90.501 90.899 90.77 90.81 90.77 90.747
46 90.949 90.949 90.818 90.099 90.995 90.949 90.818 90.93 90.936
Table 10: Test Accuracies(%) of FS Algorithms on Diabetes dataset, using XGBoost as Downstream classifier and Phi2 as the Backbone LM of ICE-SEARCH. ICES $n^{th}$ stands for the feature set with the $n^{th}$ highest validation accuracy chosen by ICE-SEARCH.
Table 10 presents the comparative performance of FS algorithms on the Diabetes dataset, utilizing XGBoost as a downstream classifier and Phi2 as the LM. It is important to note that the dataset, after balancing, comprises 436,668 samples and includes 21 indicative variables. This setup enables the examination of ICE-SEARCH’s effectiveness in processing and filtering a large volume of information.
In four out of five seeds, ICE-SEARCH $1^{st}$ identifies feature sets with higher test accuracies than those recommended by the four classical FS methods. Furthermore, other feature sets in the final pool of ICE-SEARCH also exhibit high performance. Specifically, at seeds 42 and 43, the ICE-SEARCH $2^{nd}$ discovers feature sets with test accuracies exceeding those selected by the four FS methods and ICE-SEARCH $1^{st}$; at seed 44, ICE-SEARCH $3^{rd}$ identifies the feature set with the highest test accuracy compared to those chosen by the four FS methods and ICE-SEARCH $1^{st}$.
In summary, ICE-SEARCH consistently outperforms traditional feature selection (FS) methods and is capable of identifying a greater number of high-quality feature sets from a large dataset. These findings provide empirical evidence that establishes the ICE-SEARCH as a superior FS technique compared to conventional FS methods.
4.6 Rate of Convergence:
(a) Average Top 5 Training and Validation Accuracies from 10-Fold Cross-Validation for the Stroke Dataset
(b) Average Top 5 Training and Validation Accuracies from 10-Fold Cross-Validation for the Cardiovascular Disease Dataset
(c) Average Top 5 Training and Validation Accuracies from 10-Fold Cross-Validation for the Diabetes Dataset
Figure 4: Comparative convergence analysis of XGBoost model accuracies across stroke, cardiovascular, and diabetes datasets over 10-fold cross-validation. The top row displays the convergence trends for the stroke dataset, the middle for the cardiovascular dataset, and the bottom figure for the diabetes dataset. This figure highlights the model’s performance and convergence behavior in handling different medical datasets through cross-validation.
ICE-SEARCH algorithm demonstrates fast convergence characteristics in isolating optimal sets of features. Specifically, as shown in Fig 3(a), the accuracy increases steadily until approximately Epoch three, after which the rate of improvement diminishes and the lines become parallel to the epoch axis, suggesting that the model is nearing their performance limits.
In Fig 3(b), the accuracy improves consistently over a larger number of epochs, not plateauing until around Epoch five. This suggests a slower learning process, possibly due to a more complex feature structure requiring more iterations to reach optimal performance.
In Fig 3(c), it demonstrates a rapid initial improvement in accuracy, with convergence mostly occurring by Epoch one. Post Epoch one, the accuracy levels off, indicating that the model located a solution space that did not significantly improve with further training.
In each case, convergence is indicated by the leveling out of accuracy improvements, where additional epochs do not result in significant gains.
Overall, the convergence is approximately attained by Epoch three, five, and one in the three cases. These convergence patterns observed are indicative of the swift learning capabilities of ICE-SEARCH. Specifically, the initial steep slope observed in each chart, followed by a plateau phase, highlights the efficiency of early learning. It is also worth noting that such patterns of convergence are desirable as they indicate a model’s ability to rapidly approximate the optimal function without the need for excessive training.
5 Discussion
To effectively improve the metrics of interest and search for valuable feature sets, ICE-SEARCH employs a LM for crossover and mutation operations and classical FS algorithms for initialization. We have noticed five key advantages of our proposed approach:
High generalizability: ICE-SEARCH can be readily adapted to a variety of tasks and algorithms. As evidenced by Tables (1, 4, 6), ICE-SEARCH continues to output high-performance feature subsets when integrated with different backbone LMs and downstream predictors, shedding light on the generalizability of ICE-SEARCH. Additionally, Table 10 showscases our algorithm’s ability to manage high-dimensional and high-volume datasets.
Robust against Poor Initialization: As illustrated in Tables (1, 3), ICE-SEARCH has shown robustness to ineffective initialization methods and maintained its feature selection efficiency across three distinct downstream binary classification tasks. It exhibits the capability to adapt to various initialization methods while preserving robustness against the quality of the initial population.
Adaptability: ICE-SEARCH can be easily adapted to datasets with imbalanced training and testing distributions. For example, Table 9 shows that its alternative, Decision-Randomized ICE-SEARCH maintains high performances despite differences in the distributions of training and test datasets, particularly evident in the cardiovascular disease dataset. This demonstrates textitDecision-Randomized ICE-SEARCH’s robustness to discrepancies between the distributions of training and test datasets, thereby highlighting the overall adaptability of ICE-SEARCH.
Fast Convergence: As shown in Fig 4, on all three datasets, ICE-SEARCH achieves convergence within three to four epochs, showcasing an efficient convergence rate in identifying the optimal feature sets.
Light-weight training design: ICE-SEARCH leverages the extensive world knowledge pre-trained into a LM to iteratively refine a population of feature sets. Figs 1 and 2 provide a high-level illustration of our implementtaion of ICE-SEARCH. It employs strategies such as role-playing and few-shot learning without the need for actual fine-tuning. In addition, more advanced techniques such as Tree-of-Thoughts prompting can be applied to the framework ( Yao et al. (2023)). This training approach effectively reduces the computational demands associated with training high-performance deep learning (DL) models, as highlighted by  Thompson et al. (2020).
6 Conclusion
LMs with their adeptness at encapsulating vast amounts of world knowledge and facilitating cross-domain insights, have demonstrated exceptional utility across a wide range of fields. This paper introduces ICE-SEARCH, which for the first time explores the integration of LMs with an evolutionary search framework for FS in MPA tasks.
ICE-SEARCH harnesses the crossover and mutation capabilities inherent in LMs within an evolutionary framework, significantly improving FS through the model’s comprehensive world knowledge and its adaptability to a variety of roles. Our evaluation of this methodology spans three crucial MPA tasks: stroke, cardiovascular disease, and diabetes, where ICE-SEARCH outperforms traditional FS methods in pinpointing essential features for medical applications.